OpenStudy (anonymous):

It takes Bob four hours longer to repair a car than it takes Ken. Working together they cab complete the job in 1.5 hours. How long would each of them take working alone?

hero (hero):

Actually, this one will take a bit more thought than the previous one, but the same formula still applies.

hero (hero):

B = hours Bob can repair a car alone K = hours Ken can repair a car alone x = What Bob and Ken can repair while working together The formula is: \[\frac{B \times K}{B + K} = x\]

hero (hero):

Now, it takes Bob four hours longer to repair so

hero (hero):

B = K + 4

hero (hero):

We replace B with K + 4 in the formula Also we know x = 1.5

hero (hero):

\[\frac{(K + 4) \times K}{K + 4 + K} = 1.5\]

OpenStudy (anonymous):

Ok so would I subtract 4 from 1.5? Or make it 4K * K?

hero (hero):

Do you know how to expand K(K + 4) ?

hero (hero):

You have to expand K(K + 4) for the numerator

OpenStudy (anonymous):

SO K^2 + 4K?

hero (hero):

Very good

hero (hero):

For the denominator you have to add K + K + 4

OpenStudy (anonymous):

4K^2?

hero (hero):

K + K + 4 = 2K + 4

hero (hero):

So what you have is \[\frac{K^2 + 4K}{2K + 4} = 1.5\]

hero (hero):

The afterwards, you have to cross multiply to get \[K^2 + 4K = 1.5(2K + 4)\]

hero (hero):

From there, you solve for K

OpenStudy (anonymous):

K^2+4K=3K+6

hero (hero):

Very good. Now subtract 3K from both sides. Then subtract 6 from both sides. After doing that, you should get a quadratic equation of the form ax^2 + bx + c = 0

hero (hero):

When you get that, solve for K by factoring or quadratic formula.

OpenStudy (anonymous):

K^2+K+6=0 is this right?

hero (hero):

Almost. You have to SUBTRACT 6 from both sides.

OpenStudy (anonymous):

Ok so would the factoring be. (x+3) (-2)

hero (hero):

After subtracting 6 from both sides you should have gotten K^2 + K - 6 = 0

hero (hero):

Then by factoring, you would get (K + 3)(K - 2) = 0

OpenStudy (anonymous):

Ok I got that. Ok I meant to put K. (Used to using X).

hero (hero):

When you solve for K, you must remember that time only makes sense when positive. So keep the positive result and discard the negative result.

hero (hero):

Knowing this, what is the correct value of K?

OpenStudy (anonymous):

3

hero (hero):

Not exactly. You did not solve for K

hero (hero):

To Solve for K, you have to use zero product property for (K + 3)(K - 2) = 0

hero (hero):

You solve these two equations to find possible values of K K + 3 = 0 K - 2 = 0

OpenStudy (anonymous):

-3 and 2 ?

hero (hero):

K = -3 or K = 2 But remember, K represents the amount of hours Ken can do the job alone. -3 hours makes no sense.

hero (hero):

So only K = 2 makes sense.

OpenStudy (anonymous):

Right. So Ken can make the repairs in 2 hours.

hero (hero):

Knowing this, how fast can Bob repair the car?

OpenStudy (anonymous):

6 hours.

hero (hero):

Congratulations. That is correct.

OpenStudy (anonymous):

Thanks so much! :)

hero (hero):

We can check this \[\frac{2 \times 6}{2 + 6} = 1.5\]

hero (hero):

We should get 1.5 for the left side after calculation